An elevator in on the thirtieth floor of an office building when supporting cables shears.

The elevator drops vertically to the ground where several springs absorb the impact of the

elevator. The elevator mass is 2500 kg and it is 100 m. above the ground. Determine: a)

the potential energy of the elevator before its fall, b) the velocity and kinetic energy the

instant before the impact.


1
Expert's answer
2021-11-09T10:58:22-0500

a) PE = mgh = (2500 kg)(9.81 m/s )(100 m) = 2,452,500J


b) Since we can assume there are no outside forces acting on theelevator, potential energy is fully converted into kinetic energywhen the elevator hits the springs. Therefore, KE = 2452.5 kJ

As for velocity..

KE=(1/2)mv ∴ v = √2KE/m = √2(2,452,500)/(2500) = 44.29 m/s